OpenStudy (anonymous):

how do i change y+3=6/5 x + 12/5 to a general linear form?

Parth (parthkohli):

\( \color{Black}{\Rightarrow y + 3 = \Large {6x \over 5} + {12 \over 5}}\) Since the denominators are equal, you may add both the numerators. Okay?

Parth (parthkohli):

Are you there? @dorkkk

OpenStudy (anonymous):

yes

Parth (parthkohli):

So did you get my point right there? :)

OpenStudy (anonymous):

so y+3=18/5x

Parth (parthkohli):

No! \(6x + 12\) remains as it is.

OpenStudy (anonymous):

?

Parth (parthkohli):

\( \color{Black}{\Rightarrow y + 3 = \Large{6x + 12 \over 5} }\) Do you get it till this point?

OpenStudy (anonymous):

yes